3)Find the angle between the plane 2x-6z + 5= 0 and the line passing through 2 points (1, 1, 1) and (-2, 1, 4). Hint, the angle between a line and a plane is the angle between direction vector of the line and the normal vector of the plane.
Answer #1

zu-67 +5=0. . 2u +0.4-6 7+5=0. - 1 passing through are. the directin ratios of the plane are. = 2, m=0, n. =-6 Now, the direc